If alpha and beta are the roots of the equation 3x2 -2x- 1=0.
Find 1. Alpha - beta
2. Apha cubed - beta cubed
1
Expert's answer
2019-10-28T11:28:43-0400

Roots of this equation are "1" and "-\frac{1}{3}"

1.Alpha - beta is equal "\pm(1-(-\frac{1}{3}))=\pm\frac{4}{3}"

2.Apha cubed - beta cubed is equal "\pm(1^3-(-\frac{1}{3})^3)=\pm(1+\frac{1}{27})=\pm\frac{28}{27}"

Ansswer: 1) "\pm\frac{4}{3}" , 2) "\pm\frac{28}{27}."